# Delays in Spiking Neural Networks: A State Space Model Approach

### v1

@@
Delays in Spiking Neural Networks:
A State Space Model Approach
Sanja Karilanova, Subhrakanti Dey, Ayça Özçelikkale
Department of Electrical Engineering, Uppsala University, Sweden
{Sanja.Karilanova, Subhrakanti.Dey, Ayca.Ozcelikkale}@angstrom.uu.se
S. Karilanova and A. Özçelikkale acknowledge the support of Center for Interdisciplinary Mathematics (CIM), and AI4Research, Uppsala University.The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
I Abstract
@@
Spiking neural networks (SNNs) are biologically inspired, event-driven models that are suitable for processing temporal data and offer energy-efficient computation when implemented on neuromorphic hardware. In SNNs, richer neuronal dynamic allows capturing more complex temporal dependencies, with delays playing a crucial role by allowing past inputs to directly influence present spiking behavior. We propose a general framework for incorporating delays into SNNs through additional state variables. The proposed mechanism enables each neuron to access a finite temporal input history. The framework is agnostic to neuron models and hence can be seamlessly integrated into standard spiking neuron models such as LIF and adLIF. We analyze how the duration of the delays and the learnable parameters associated with them affect the performance. We investigate the trade-offs in the network architecture due to additional state variables introduced by the delay mechanism. Experiments on the Spiking Heidelberg Digits (SHD) dataset show that the proposed mechanism matches the performance of existing delay-based SNNs while remaining computationally efficient. Moreover, the results illustrate that the incorporation of delays may substantially improve performance in smaller networks. Code available at https://github.com/sannkka/Modeling-State-Delays-in-SNNs.
II Introduction
@@
Spiking Neural Networks (SNNs) [MAASS19971659] have emerged as a more biologically plausible class of neural models compared to conventional artificial neural networks (ANNs), offering promising advantages in temporal processing and energy efficiency, particularly when deployed on neuromorphic hardware [davies_advancing_2021]. SNNs encode information in the timing of discrete spikes, allowing them to capture temporal structure and making them well suited for modeling time-dependent signals such as audio [bittar2022surrogate], sensory [10.3389/fnins.2022.951164], or visual patterns [Gallego2022]. Despite these advantages, achieving performance comparable to that of ANNs remains an open challenge, in part due to the limited temporal expressivity of current SNN models.
@@
In biological neural systems, temporal delays play a critical computational role. Delays that enable neurons to integrate information from multiple past moments support tasks such as hearing [carr1993processing] and brain activity [carr1990circuit]. Inspired by these findings, numerous works have explored incorporating delays into computational SNNs models. Theoretical analyses have shown that an SNN with $k$ adjustable delays can compute a much richer class of functions than a threshold circuit with $k$ adjustable weights [MAASS199926]. Empirical studies have also demonstrated improved performance when optimizing spike transmission delays along with synaptic weights or membrane time constants [deckers2024co, sun2025exploitingheterogeneousdelaysefficient]. These results suggest that delays can enrich the information processing capabilities of SNNs and their internal representation. However, despite these promising results, practical methods for introducing structured, scalable, and differentiable forms of delay into modern deep SNNs remains an ongoing research problem.
@@
In this work, we propose a method based on State Space Models (SSMs) for equipping SNNs with delays. The method uses a time-shift based state transition, which enables each neuron to access both its current input and the most recent history of past input values.The formulation is general and can be incorporated into arbitrary spiking neuron models. Conceptually, the framework bridges ideas from SSMs and SNNs, allowing SNNs to capture a fixed number of past inputs in a structured and interpretable manner.
@@
  • We propose a method for incorporating delays within the state dynamics through additional delay variables, enabling neurons to access a finite temporal history of inputs.
@@
  • The proposed formulation is general and easily integrable into existing spiking neuron models, including LIF, adLIF, and RadLIF.
@@
  • We conduct a systematic analysis of delay order, i.e. the number of allowed past inputs, and their associated learnable parameters, and study the trade-offs between the number of neurons in the network and the delay order.
@@
  • The proposed framework introduces either no additional trainable parameters (i.e when the delay parameters are not trained) or a number of trainable parameters comparable to the other SNN delay models.
@@
The proposed models achieve performance comparable to existing delay-based SNN approaches on the Spiking Heidelberg Digits (SHD) dataset. Moreover, the observed trade-offs highlight promising directions for future exploration, particularly showing that incorporating delays can substantially improve accuracy in smaller SNNs where the neuron count limits temporal capacity.
Figure 1: Illustration of different approaches of modeling delays. Our proposed delays are modeled within the state dynamics of the neuron (green). Figure is inspired by [DelGrad, Fig. 3(a)].
III Related Work
@@
In the SNN literature, three main types of delays are commonly modeled: axonal delays, applied to the output of a neuron [shrestha2018slayer, 10.3389/fnins.2023.1275944, sun2025exploitingheterogeneousdelaysefficient]; dendritic delays, applied to its input [diaz2016efficient]; and synaptic delays, specific to each connection between neuron pairs [10.1145/3589737.3606009, deckers2024co, hammouamri2023learning, dominijanni2025extendingspiketimingdependentplasticity, meszaros2024learning]. In contrast to these formulations, we propose an SSM based approach where delays are modeled within the state dynamics of the neuron. See Figure 1 for a schematic illustration.
The most common ways to model delays in SNNs is to learn a single delay value applied to each spike [shrestha2018slayer], or to employ temporal convolutions [hammouamri2023learning]. Several works explore the role of delays in SNNs by jointly training weights and delays [deckers2024co], learning only delays [10.1145/3589737.3606009], or showing that heterogeneous delays can reduce weight-precision requirements [sun2025exploitingheterogeneousdelaysefficient]. Others relate delays to temporal parameters such as time constants [10.1371/journal.pcbi.1012673] or study their role in the formation of temporal hierarchies between layers [moro2024roletemporalhierarchyspiking]. In these works, various learning algorithms have been used, such as variations of backpropagation through time (BPTT) [hammouamri2023learning, shrestha2018slayer, queant2025delreclearningdelaysrecurrent, meszaros2024learning], biologically inspired spike-timing-dependent plasticity (STDP) [dominijanni2025extendingspiketimingdependentplasticity], and exact spike-time gradients [DelGrad, meszaros2025efficient]. Our formulation represents delays as state variables evolving over time. This structure remains compatible with BPTT, enabling training with the popular surrogate-gradient methods [neftci2019surrogate] used to train SNNs.
@@
Delays have long been studied in fields such as signal processing and control theory. In discrete-time dynamical systems, they are typically modeled through delay-difference or state-space formulations, where past states are explicitly included to capture finite memory effects [RICHARD20031667, SenameOlivier2005Sots]. Recently, modern deep SSMs, which stack multiple state-space layers combined with nonlinearities, have demonstrated strong performance in sequence modeling tasks [gu2022efficientlymodelinglongsequences, Gupta_DSS]. Even more recently, parameterization has been introduced in deep SSMs, effectively capturing delayed state dependencies [fu2022hungry, zhang2023effectively].
We build on this formulation to represent controllable delays within spiking neural network models. A common approach in modern SNNs embeds spiking dynamics within SSMs, where at each time step the output of the deep SSM layer is passed through a firing threshold function [stan2023learning, bal2024rethinkingspikingneuralnetworks, karilanova2025statespacemodelinspiredmultipleinput, karilanova_reset, du2024spiking]. We continue in this direction by utilizing the SSM perspective on delay modeling to spiking neural architectures, enabling SNNs to incorporate structured temporal dependencies through the same principled state-space formulation.
IV Methods
IV-A Preliminaries
IV-A1 General Spiking Neuron Model
@@
A single general feed-forward discrete-time $n$-state dimensional spiking neuron can be defined as [karilanova2025zeroshottemporalresolutiondomain, karilanova2025statespacemodelinspiredmultipleinput, karilanova_reset]:
@@
The neuron model in (1) generalizes several commonly used formulations, including the Leaky Integrate-and-Fire (LIF) and adaptive LIF (adLIF) neurons, see [karilanova2025zeroshottemporalresolutiondomain, karilanova2025statespacemodelinspiredmultipleinput, karilanova_reset] for illustration of such representations.
Spiking neural networks are built by connecting neurons such that the output spikes from the previous (pre-synaptic, l1l-1) layer serve as inputs to the next (post-synaptic, $l$) layer. Let $L_{\text{pre}}$ and $L_{\text{post}}$ be the number of neurons in the pre and post synaptic layer. Let $\bm{W}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{pre}}}$ be the weight matrix connecting the two layers, and $\bm{V}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{post}}}$ be the weight matrix associated with the recurrent connections between neurons in a layer. Then $\bm{i}[t]=\bm{W}\times\bm{s}^{l-1}[t]+\bm{V}\times\bm{s}^{l}[t]$, where $\bm{s}^{l-1}[t]$ and $\bm{s}^{l}[t]$ are a vector of all the spikes from the previous and current layer, respectively.
IV-A2 State Space Model with Time Shift State Transition Matrix
A discrete-time time-invariant linear SSM can be written as [ljung_SI]:
@@ IV-A2 State Space Model with Time Shift State Transition Matrix
The lower shift matrix is a structured form of the state-transition matrix with nonzero elements only on the subdiagonal. Let $\bm{A}_{d}$ be the lower shift matrix:
$\bm{A}_{d}$ $=\begin{bmatrix}0&0&\ldots&0&0\\ a_{1}&0&\ldots&0&0\\ 0&a_{2}&\ldots&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&a_{n_{d}-1}&0\end{bmatrix}\text{i.e.,}$ (3a)
$[\bm{A}_{d}]_{i,j}$ $=\begin{cases}a_{j}&\text{ for }i-1=j\\ 0&\text{ otherwise }\end{cases}$ (3b)
where $\bm{A}_{d}]_{i,j}$ represents the $i$th row $j$th column element of the matrix $\bm{A}_{d}$. Let
$\bm{B}_{d}=[1,0,\ldots,0]^{T}.$ (4)
Then for the SSM state variable $\bm{v}_{d}$ defined in (2) and initial state $\bm{v}_{d}[0]=\bm{0}^{n_{d}\times 1}$ we have:
$\bm{v}_{d}[1]$ $=[i_{d}[1],\,0,\,0,\,...,\,0]^{T}$
$\bm{v}_{d}[2]$ $=[i_{d}[2],\,a_{1}i_{d}[1],\,0,\,...,\,0]^{T}$
$\bm{v}_{d}[3]$ $=[i_{d}[3],\,a_{1}i_{d}[2],\,a_{2}a_{1}i_{d}[1],\,...,\,0]^{T}$
$\,\,\,\vdots$
$\bm{v}_{d}[t]$ $=[i_{d}[t],\,a_{1}i_{d}[t-1],\,a_{2}a_{1}i_{d}[t-2],\,$...,\,
$\hskip 85.35826pt(\prod^{n-1}_{k=1}a_{k})i_{d}[t-n_{d}+1]]^{T}.$
Hence, at time $t$ the state variable $\bm{v}_{d}$ is a vector with entries equal to the last $n$ inputs to the system parametrized by the parameters $a_{i}$, where $n$ is the state dimension.
Note that although the $a_{i}$ parameters are free in general, in this paper we set $a_{i}=1$ for all $i$. Under this choice, the delay component $\bm{v}_{d}$ reduces to a simple temporal memory of the input, i.e.,
$\bm{v}_{d}[t]=[i_{d}[t],i_{d}[t-1],i_{d}[t-2],...,i_{d}[t-n_{d}+1]]^{T}.$ (5)
IV-B Proposed Model
Let $\bm{v}_{s}$ be the state variable of the general spiking neuron in (1) with parameter matrices $\bm{A}_{s},\bm{B}_{s},\bm{C}_{s}$ and state dimension $n_{s}$. Let $\bm{v}_{d}$ be the state variable of the time-shift SSM in Section IV-A2 with parameter matrices $\bm{A}_{d},\bm{B}_{d},\bm{C}_{d}$ and state dimension $n_{d}$. Our proposed spiking neuron model with state-delays is defined as:
@@ IV-B Proposed Model
$s[t]$ $=f_{\bf{\Theta}}(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])$
$=\begin{cases}1\text{ if }(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])\in\bf{\Theta}\\ 0\text{ otherwise}\end{cases}$ (6c)
Here, $\bm{A}_{sd}\in\mathbb{R}^{n_{s}\times n_{d}}$ gives the parameters that delayed inputs in the state $\bm{v}_{d}$ are multiplied with before they are added to the neuron state $\bm{v}_{s}$.
The above equations present a spiking neuron with additional access to temporal memory through the state variable $\bm{v}_{d}[t]$. The state dimension $n_{d}$ defines the maximum order of the delays provided to the spiking neuron i.e. the number of directly accessible past quantities. The input $i_{s}$ to $\bm{v}_{s}$, as defined in Section IV-A1, represents the input spikes to the neuron. The input $i_{d}$ to $\bm{v}_{d}$ represents the input stored in the temporal memory. See Table I for an overview of the notation. Examples of the proposed models with LIF and adLIF neuron models are provided in Sections IV-B1 and IV-B2, respectively. A high-level visualization of the proposed model is presented in Figure 2.
TABLE I: Table of notation.

### v2

@@
Abstract
@@
Spiking neural networks (SNNs) are biologically inspired, event-driven models suited for temporal data processing and energy-efficient neuromorphic computing. In SNNs, richer neuronal dynamic allows capturing more complex temporal dependencies, with delays playing a crucial role by allowing past inputs to directly influence present spiking behavior. We propose a general framework for incorporating delays into SNNs through additional state variables. The proposed mechanism enables each neuron to access a finite temporal input history. The framework is agnostic to neuron models and hence can be seamlessly integrated into standard spiking neuron models such as Leaky Integrate-and-Fire (LIF) and Adaptive LIF (adLIF). We analyze how the duration of the delays and the learnable parameters associated with them affect the performance. We investigate the trade-offs in the network architecture due to additional state variables introduced by the delay mechanism. Experiments on the Spiking Heidelberg Digits (SHD) dataset show that the proposed mechanism matches existing delay-based SNNs in performance while remaining computationally efficient, with particular gains in smaller networks. Code available at https://github.com/sannkka/Modeling-State-Delays-in-SNNs.
1 Introduction
@@
Spiking Neural Networks (SNNs) have emerged as a more biologically plausible class of neural models compared to conventional artificial neural networks (ANNs), offering promising advantages in temporal processing and energy efficiency, particularly when deployed on neuromorphic hardware [6]. SNNs encode information in the timing of discrete spikes, allowing them to capture temporal structure and making them well suited for modeling time-dependent signals such as audio [1], or visual patterns [13]. Despite these advantages, achieving performance comparable to that of ANNs remains an open challenge, in part due to the limited temporal expressivity of current SNN models.
@@
In biological neural systems, temporal delays play a critical computational role. Delays that enable neurons to integrate information from multiple past moments support tasks such as hearing [3] and brain activity [4]. Inspired by these findings, numerous works have explored incorporating delays into computational SNNs models. Theoretical analyses have shown that an SNN with $k$ adjustable delays can compute a much richer class of functions than a threshold circuit with $k$ adjustable weights [24]. Empirical studies have also demonstrated improved performance when optimizing spike transmission delays along with synaptic weights or membrane time constants [7, 32]. These results suggest that delays can enrich the information processing capabilities of SNNs and their internal representation. However, despite these promising results, practical methods for introducing structured, scalable, and differentiable forms of delay into modern deep SNNs remains an ongoing research problem.
@@
In this work, we propose a method based on State Space Models (SSMs) for equipping SNNs with delays. The method uses a time-shift based state transition, which enables each neuron to access both its current input and the most recent history of past input values. The formulation is general and can be incorporated into arbitrary spiking neuron models, such as Leaky Integrate-and-Fire (LIF) and Adaptive LIF (adLIF) and their recurrent variant Recurrent LIF (RLIF) and Recurrent adLIF (RadLIF) [1]. Conceptually, the framework bridges ideas from SSMs and SNNs, allowing SNNs to capture a fixed number of past inputs in a structured and interpretable manner.
@@
  • We propose incorporating delays within the neuron state dynamics via additional delay variables, enabling access to a finite temporal history of inputs.
@@
  • The formulation is general and easily integrable into existing spiking neuron models, including LIF, RLIF, adLIF, and RadLIF.
@@
  • We systematically analyze delay order (number of accessible past inputs) and associated learnable parameters, studying trade-offs between network size and delay order.
@@
  • The framework introduces either no additional trainable parameters (non-trainable case) or a parameter count comparable to other SNN delay models.
@@
The proposed models achieve performance comparable to existing delay-based SNN approaches on the Spiking Heidelberg Digits (SHD) dataset. The observed trade-offs highlight promising directions for future exploration, particularly showing that incorporating delays can substantially improve accuracy in smaller SNNs where the neuron count limits temporal capacity.
2 Related Work
@@
In the SNN literature, three main types of delays are commonly modeled: axonal delays, applied to the output of a neuron [31, 33, 32]; dendritic delays, applied to its input [8]; and synaptic delays, specific to each connection between neuron pairs [15, 7, 17, 9]. In contrast to these formulations, we propose an SSM based approach where delays are modeled within the state dynamics of the neuron. See Figure 1 for a schematic illustration.
A typical approach for modeling delays in SNNs is through explicit temporal convolutions applied to spike trains [17, eqn. 6–8], [33, eqn. 4], [31, eqn. 3], where a kernel with a relatively small number of nonzero weights is convolved with the spike train, and the relative positions of the non-zero weights encode the delay mechanism. A related approach, and the one closest to our proposed method, is to incorporate temporal buffer directly into the neuron state as an explicit finite memory of past inputs [34]. Given that linear state-space models may be interpreted as temporal convolutions [12], this type of approaches can be seen as an extra incorporation of delay into temporal convolution, see also (3). An alternative line of work models delays as continuous-valued spike-time shifts [14, 25], which permits exact gradient methods but is less compatible with discrete surrogate-gradient training pipelines.
Figure 1: Delay modeling approaches. Proposed method: delays within neuron state dynamics (green). Figure inspired by [14, Fig. 3(a)].
The role of delays in SNNs has been studied from various perspectives including jointly training weights and delays [7], learning only delays [15], exploring heterogeneous delays and their relation to weight-precision requirements [32], and investigating their role in temporal hierarchies across layers [26]. Learning algorithms for delays range from BPTT [17, 31] and STDP [9] to exact spike-time gradients [14, 25]. Our formulation represents delays as state variables, making it compatible with BPTT and the widely used surrogate-gradient method for training SNNs [27].
@@
Delays have long been studied in fields such as signal processing and control theory. In discrete-time dynamical systems, they are typically modeled through delay-difference and state-space formulations [28, 30]. Modern deep SSMs, which stack multiple state-space layers combined with nonlinearities, have demonstrated strong performance in sequence modeling tasks [16]. Subsequent work introduced parameterization in deep SSMs, which effectively captures delayed state dependencies [11, 36]. In parallel, hybrid approaches have emerged that bridge SNNs and SSMs, combining properties and capabilities from both models [35, 20, 19, 10, 21, 32]. We build on this line of work and adopt the SSM perspective on delay modeling within spiking neural architectures. This allows SNNs to incorporate structured temporal dependencies through a principled state-space formulation.
3 Methods
3.1 Preliminaries
3.1.1 General Spiking Neuron Model
@@
A single general feed-forward discrete-time $n$-state dimensional spiking neuron can be defined as [21]:
@@
The model in (1) generalizes commonly used formulations including including the Leaky Integrate-and-Fire (LIF) and adaptive LIF (adLIF) neurons and their recurrent variants Recurrent LIF (RLIF) and Recurrent adLIF (RadLIF) [21].
In SNNs, output spikes from the pre-synaptic layer (l1l-1) serve as inputs to the post-synaptic layer ($l$). Let $\bm{W}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{pre}}}$ and $\bm{V}\in\mathbb{R}^{L_{\text{post}}\times L_{\text{post}}}$ be the feedforward and recurrent weight matrices, with $L_{\text{pre}}$, $L_{\text{post}}$ the layer sizes. The input current is $\bm{i}[t]=\bm{W}\times\bm{s}^{l-1}[t]+\bm{V}\times\bm{s}^{l}[t]$, where $\bm{s}^{l-1}[t]$ and $\bm{s}^{l}[t]$ are spike vectors from the previous and current layer.
3.1.2 Linear State Space Model
A discrete-time time-invariant linear SSM can be written as [16]:
@@ 3.1.2 Linear State Space Model
3.1.3 Temporal Convolution Representation of Linear SSM
A key property of linear SSMs is that it can be expressed as a causal temporal convolution. Under the initial condition $v_{d}[t]=0$, for $t\leq 0$, unrolling the recurrence in (2a) yields
$\bm{v}_{d}[t]=\sum_{k=0}^{t-1}\bm{A}_{d}^{k}\bm{B}_{d}\,i_{d}[t-k].$ (3)
The state variable at time $t$ is therefore a convolution of the input with the kernel $d[t]=\bm{A}_{d}^{t}\bm{B}_{d}$ i.e. $\bm{v}_{d}[t]=d[t]*i_{d}[t]$, where * denotes convolution.
3.1.4 Linear SSM with Time Shift State Transition Matrix
The lower shift matrix is a structured matrix whose nonzero entries lie only on the subdiagonal. Let $\bm{A}_{d}$ in (2) denote such a matrix i.e. let
$\bm{A}_{d}=\begin{bmatrix}0&0&\ldots&0&0\\ a_{1}&0&\ldots&0&0\\ 0&a_{2}&\ldots&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&a_{n_{d}-1}&0\end{bmatrix}\text{ i.e., }[\bm{A}_{d}]_{i,j}=\begin{cases}a_{j}&\text{ for }i-1=j\\ 0&\text{ otherwise }\end{cases}$ (4a)
where $\bm{A}_{d}]_{i,j}$ denotes the $(i,j)$-th entry of $\bm{A}_{d}$, and let $\bm{B}_{d}=[1,0,\ldots,0]^{T}$. Although the parameters $a_{i}$ are free in general, in this paper we set $a_{i}=1$ for all $i$. Under this choice, and following (3), $\bm{v}_{d}$ reduces to a simple temporal memory of the last $n$ inputs (a shift register), i.e.,
$\bm{v}_{d}[t]=[i_{d}[t],i_{d}[t-1],i_{d}[t-2],\ldots,i_{d}[t-n_{d}+1]]^{T},$ (5)
see Appendix 0.A.2 for a demonstration.
3.2 Proposed Model
Let $\bm{v}_{s}$ be the state variable of the general spiking neuron in (1) with parameter matrices $\bm{A}_{s},\bm{B}_{s},\bm{C}_{s}$ and state dimension $n_{s}$. Let $\bm{v}_{d}$ be the state variable of the time-shift SSM in Section 3.1.4 with parameter matrices $\bm{A}_{d},\bm{B}_{d},\bm{C}_{d}$ and state dimension $n_{d}$. Our proposed spiking neuron model with state-delays is defined as:
@@ 3.2 Proposed Model
$s[t]$ $=f_{\bf{\Theta}}(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])=\begin{cases}1\text{ if }(\bm{C}_{s}\bm{v}_{s}[t]+\bm{C}_{d}\bm{v}_{d}[t])\in\bf{\Theta}\\ 0\text{ otherwise}\end{cases}$ (6c)
Here, $\bm{A}_{sd}\in\mathbb{R}^{n_{s}\times n_{d}}$ denotes the parameters that delayed inputs in the state $\bm{v}_{d}$ are multiplied with before they are added to the neuron state $\bm{v}_{s}$.
The above equations present a spiking neuron with additional access to temporal memory through the state variable $\bm{v}_{d}[t]$. The state dimension $n_{d}$ defines the maximum order of the delays provided to the spiking neuron i.e. the number of directly accessible past quantities. The input $i_{s}$ to $\bm{v}_{s}$, as defined in Section 3.1.1, represents the input spikes to the neuron. The input $i_{d}$ to $\bm{v}_{d}$ represents the input stored in the temporal memory. See Table 1 for an overview of the notation. Examples of the proposed models with LIF, RLIF, adLIF and RadLIF neuron models are provided in Appendix 0.A.1. A high-level visualization of the proposed model is presented in Figure 2.
Table 1: Table of notation.